# Emergent Weyl-like points in periodically modulated systems and Floquet engineering

### v3

@@ Abstract
We investigate a three-dimensional (3D) topological phase resembling a Weyl semimetal, modulated by a periodic potential and engineered through Floquet dynamics. This system is constructed by stacking two-dimensional Chern insulators and hosts Weyl-like points defined in the parameter space $(k_{x},k_{y},z)$, distinct from conventional Weyl points in momentum space $(k_{x},k_{y},k_{z})$. The Weyl-semimetal-like phase exhibits characteristics akin to those of Weyl semimetals, including linear dispersion near the Weyl-like points, nontrivial bulk topology, and the presence of Fermi arcs connecting the Weyl-like points. Unlike traditional Weyl semimetals, these features manifest in real space rather than momentum space. Furthermore, we calculate the local density of states, the layer Hall conductance, and the total 3D Hall conductivity, demonstrating that the Weyl-semimetal-like phase remains stable under weak and moderate interlayer couplings. Finally, we compute the Floquet Chern number and show that the positions of the Weyl-like points can be tuned via high-frequency laser pumping.
@@ I Introduction
Stacks of two-dimensional (2D) Chern insulator phases have proven to be a versatile platform for realizing extraordinary topological phases, attracting considerable attention in recent years Mogi et al. (2017a, b); Xiao et al. (2018); Mogi et al. (2022); Qin et al. (2023); Chen and Shen (2024); Chang et al. (2013); Yoshimi et al. (2015); Shan et al. (2010); Lu et al. (2010); Li et al. (2010); Lu et al. (2013); Chen et al. (2019); Sun et al. (2020); Liu et al. (2010); Dabiri et al. (2021); Dabiri and Cheraghchi (2021); Sun et al. (2023); Qin et al. (2022a); Zhu et al. (2023). For instance, the well-known Weyl semimetal phase can be interpreted as a stack of 2D Chern insulators with orbital-dependent interlayer tunneling Yang et al. (2011); Chen et al. (2015, 2018a, 2018b); Li et al. (2021). Similarly, phases such as the axion insulator and the hinged quantum spin Hall effect arise from alternating stacks of Chern insulators with opposite Chern numbers Chen et al. (2021, 2024a, 2023a, 2023b). Moreover, higher-Chern-number quantum anomalous Hall insulators can be constructed by alternating stacks of Chern insulators and normal insulators Zhao et al. (2022); Baba et al. (2022); Li et al. (2022, 2023).
@@
Zhang et al. recently proposed and studied a 2D phase modulated by a periodic potential Zhang et al. (2023). Unlike conventional Chern insulators, where the spectral flow of edge modes is momentum-dependent, their system exhibits edge modes whose spectral flow depends on spatial position. They also suggested that such a periodic potential could be experimentally realized via charge density wave effects Huang et al. (2021); Qin et al. (2020). In a related development, Chen et al. explored a 2D higher-order topological insulator and demonstrated that the energies of the corner modes can be tuned by a periodic spin-orbital potential Chen et al. (2024b).
@@
Off-resonant light serves as a powerful tool for inducing Floquet Weyl semimetals Wang et al. (2014); Hübener et al. (2017). Building on Floquet theory Oka and Aoki (2009); Zhan et al. (2024); Qin et al. (2023, 2022b, 2022a, 2024); Dabiri et al. (2021); Dabiri and Cheraghchi (2021); Dabiri et al. (2022); Cheraghchi and Askarpour (2023); Pervishko et al. (2018); Calvo et al. (2015); Seshadri and Sen (2022); Seshadri et al. (2019); Seshadri (2023); Lee et al. (2018); Magnus (1954); Blanes et al. (2009); Bukov et al. (2015); Eckardt and Anisimovas (2015); Chen et al. (2018c, d); Du et al. (2022); Wang et al. (2023), it would be intriguing to investigate the experimental implications of realistic periodic driving frequencies. This is particularly relevant in light of recent developments, such as the photoinduced anomalous Hall effect in graphene McIver et al. (2020) and its corresponding microscopic theory Sato et al. (2019), the experimental observation of Floquet states in graphene Merboldt et al. (2025), and the direct detection of Floquet-Bloch states in monolayer graphene Choi et al. (2025), WSe2 Aeschlimann et al. (2021), and the 3D topological insulator Bi2Se3 Wang et al. (2013). Additionally, experimental studies on Floquet transport effects, including anomalous linear transport in light-driven 3D Dirac electrons in bismuth Hirai et al. (2023) and nonperturbative nonlinear transport in a magnetic Floquet-Weyl semimetal Day et al. (2024), further highlight the potential of this approach.
@@
In this work, we propose and study a 3D topological phase resembling a Weyl semimetal, modulated by a periodic potential and realized through a simple stack of 2D Chern insulators. The system features Weyl-like points defined in the parameter space $(k_{x},k_{y},z)$, distinct from the conventional Weyl points in momentum space $(k_{x},k_{y},k_{z})$. This Weyl-semimetal-like phase exhibits key properties of a Weyl semimetal, including linear dispersion near the Weyl-like points, nontrivial bulk topology, and the emergence of Fermi arc connecting these points. However, these behaviors are observed in real space, rather than momentum space, as in traditional Weyl semimetals. We further compute the local density of states, the layer Hall conductance, and the total 3D Hall conductivity, demonstrating that the Weyl-semimetal-like phase remains robust under weak and moderate interlayer couplings. Additionally, we evaluate the Floquet Chern number under high-frequency periodic driving and show that the positions of the Weyl-like nodes can be effectively tuned by high-frequency laser fields.
@@
The structure of this paper is as follows. In Section II, we introduce the Weyl-like model Hamiltonian with a periodic potential along the $z$-direction, alongside the conventional Weyl semimetal model, and discuss the duality between these two models. Section III presents calculations of the local density of states, the layer Hall conductances, and the total 3D Hall conductivity in the presence of interlayer hopping for the Weyl-like model. In Section IV, we derive the Floquet effective Hamiltonians for both models in Section II under high-frequency laser driving. Section V details the analysis of Floquet Weyl nodes, Floquet Chern numbers, the Floquet energy spectra, and Floquet Fermi arcs for both models. A brief discussion of possible experimental realizations of the Weyl-like model is given in Section VI. Finally, Section VII summarizes the main results of the paper.
@@ II Model
In this section, we will introduce the Weyl-like model and the Weyl semimetal model.
@@ II.1 Weyl-like model
We begin by introducing a two-band model Hamiltonian that describes a stack of 2D Chern insulators subjected to a periodic potential along the $z$-direction. This model is defined in the $(k_{x},k_{y},z)$ space as
[…]
where $h_{\text{2D}}({\bf k}_{||})\!=\!m_{z}\sigma_{z}\!+\!m_{0}[2\!-\!\cos(k_{x}a)\!-\!\cos(k_{y}a)]\sigma_{z}\!+\!t_{x}\sin(k_{x}a)\sigma_{x}\!+\!t_{y}\sin(k_{y}a)\sigma_{y}$ represents the Hamiltonian of a 2D Chern insulator when $m_{z}\times m_{0}<0$ Shen (2017), and ${\bf k}_{||}\!=\!(k_{x},k_{y})$. Here, $a$ is the lattice constant, and $\sigma_{i}$ $(i\!=\!x,y,z)$ are Pauli matrices acting on the orbital space. The spatially periodic potential is given by $V(z)\!=\!V_{z}\cos(Q_{z}z)$, where $V_{z}$ is the potential strength, $z\!=\!j_{z}a$ denotes the position of the $j_{z}$-th layer along the $z$-direction, and $\lambda_{z}\!=\!2\pi/Q_{z}$ is the wavelength of the modulation.
@@
Figure 1: Duality between the decoupled Weyl-like model (left column) and the Weyl semimetal model (right column). The two green vertical dashed lines indicate the positions of the two Weyl-like (or Weyl) points. (a1) Schematic of the decoupled Weyl-like model in real space, featuring a spatially periodic potential $V(z)\!=\!V_{z}\cos(Q_{z}z)$, where $Q_{z}\!=\!2\pi/\lambda_{z}$. (b1) Weyl-like cones in the $(k_{x},k_{y},z)$ space at two critical planes, $(k_{x},k_{y},z_{c-})$ and $(k_{x},k_{y},z_{c+})$, marking topological transitions. (c1) Chern number of each layer [Eq. (3)] as a function of layer index $z$ for the decoupled Weyl-like model with $V_{z}\!=\!1$ eV, $\lambda_{z}\!=\!L_{z}\!=\!31a$, and the lattice constant $a\!=\!1$ nm. (d1) Energy spectrum [Eq. (6)] (blue solid curves) of the Weyl-like model at $k_{x}\!=\!k_{y}\!=\!0$. (e1) Energy spectrum (blue solid curves) and Fermi-like arc (blue horizontal line for $E\!=\!0$) for the Weyl-like model under open boundary conditions (OBCs) along the $y$-direction, with $k_{x}\!=\!0$ and system size $L_{y}\!=\!31a$. (a2) Schematic of the Weyl semimetal model in real space, featuring a momentum-dependent potential $U(k_{z})\!=\!t_{z}\cos(k_{z}a)$. (b2) Weyl cones in momentum space $(k_{x},k_{y},k_{z})$ at two critical planes $(k_{x},k_{y},-k_{c})$ and $(k_{x},k_{y},+k_{c})$, indicating the topological features of the model. (c2) Chern number [Eq. (II.2)] as a function of $k_{z}$ for the Weyl semimetal model with $t_{z}\!=\!1$ eV. (d2) Energy spectrum [Eq. (13)] (black solid curves) of the Weyl semimetal model at $k_{x}\!=\!k_{y}\!=\!0$. (e2) Energy spectrum (black solid curves) and Fermi arc (black horizontal line at $E\!=\!0$) for the Weyl semimetal model under OBCs along the $y$-direction, with $k_{x}\!=\!0$ and system size $L_{y}\!=\!31a$. Other parameters used: $m_{z}\!=\!0.5$ eV, $m_{0}\!=\!1$ eV, $t_{x}\!=\!t_{y}\!=\!t_{0}\!=\!1$ eV, and $L_{z}\!=\!31a$.
As illustrated in Fig. 1(a1), the system described by the Weyl-like Hamiltonian in Eq. (1) can be interpreted as an array of decoupled 2D layers, each with a spatially dependent effective mass term $\tilde{m}_{z}(z)\!=\!m_{z}\!+\!V(z)$. For the calculations presented in this work, we adopt the fixed parameters: $m_{z}\!=\!0.5$ eV, $m_{0}\!=\!1$ eV, $t_{x}\!=\!t_{y}\!=\!t_{0}\!=\!1$ eV, $V_{z}\!=\!1$ eV, and $\lambda_{z}\!=\!31a$ with $a\!=\!1$ nm. The detailed matrix representation of the real-space Hamiltonian with the periodic potential along the $z$-direction – assuming open boundary conditions (OBCs) along the $y$-direction and periodic boundary conditions (PBCs) along the $x$-direction – is provided in Section SI of the Supplemental Material Sup .
@@
Since the layers are decoupled from each other, we can define a $z$-dependent Chern number of each layer for the $m$-th occupied band as Fukui et al. (2005); Sun et al. (2020); Otrokov et al. (2019); Gao et al. (2021); Chen et al. (2024a); Mong et al. (2010); Essin et al. (2009); Wang et al. (2015); Varnava and Vanderbilt (2018); Chen and Shen (2024); Fu et al. (2021); Price et al. (2015); Zhang et al. (2019)
@@
The $z$-dependent Chern number of each layer is illustrated in Fig. 1(c1). Figures 1(a1) and 1(b1) show that each layer behaves as a Chern insulator with Chern number $|C|\!=\!1$ for $z_{c-}<z<z_{c+}$, and as a normal insulator with $C\!=\!0$ outside this region (i.e., for $z<z_{c-}$ and $z>z_{c+}$). To determine the critical values $z_{c\pm}$, we evaluate the eigenenergies of the Weyl-like Hamiltonian (1) at the high-symmetry point $k_{x}\!=\!k_{y}\!=\!0$, obtaining from Eq. (II.1)
@@
Furthermore, the critical values $z_{c\pm}$ correspond to the gap-closing points where $E_{\pm}\!=\!\pm\tilde{m}_{z}(z)\!=\!\pm\left[m_{z}\!+\!V_{z}\cos(Q_{z}z)\right]\!=\!0$, leading to
@@
under the condition $-1\leqslant m_{z}/V_{z}\leqslant 1$.
Figure 1(d1) shows the energy spectrum as a function of position $z$ at $k_{x}\!=\!k_{y}\!=\!0$. A clear band inversion occurs near $z\!=\!z_{c\pm}$, highlighting the spectral flow characteristic of the layered topological structure. Additionally, Fig. 1(e1) reveals the presence of a Fermi arc, further demonstrating the system’s topological nature.
Moreover, the total 3D Hall conductivity for the $m$-th occupied band is given by Burkov and Balents (2011)
[…]
where $\sigma_{H}^{2D}(z)\!=\!(e^{2}/h)C_{m}(z)$ is the Hall conductance of each layer associated with the $m$-th occupied band, $C_{m}(z)$ is defined in Eq. (3), and $L_{z}$ denotes the system size along the $z$-direction. Using the critical values $z_{c\pm}$ obtained in Eq. (7), we can simplify the 3D Hall conductivity expression as
[…]

### v4

@@ Abstract
We investigate a three-dimensional (3D) topological phase resembling a Weyl semimetal, modulated by a periodic potential and engineered through Floquet dynamics. This system is constructed by stacking two-dimensional Chern insulators and hosts Weyl-like points defined in the parameter space $(k_{x},k_{y},z)$, distinct from conventional Weyl points in momentum space $(k_{x},k_{y},k_{z})$. The Weyl-semimetal-like phase exhibits characteristics akin to those of Weyl semimetals, including linear dispersion near the Weyl-like points, nontrivial bulk topology, the presence of Fermi arcs connecting the Weyl-like points, and the Berry monopoles. Unlike traditional Weyl semimetals, these features manifest in real space rather than momentum space. Furthermore, we calculate the local density of states, the layer Hall conductance, and the total 3D Hall conductivity, demonstrating that the Weyl-semimetal-like phase remains stable under weak and moderate interlayer couplings. The influence of disorder is also examined: Beyond a critical disorder strength, the Weyl-like points destabilize and the topological phase collapses. Moreover, by computing the Floquet Chern number, we demonstrate that the locations of the Weyl-like points can be tuned via high-frequency laser pumping. Finally, we show that both type-I and II Weyl-like behaviors can arise in a tilted Weyl-like model.
@@ I Introduction
Stacks of two-dimensional (2D) Chern insulator phases have proven to be a versatile platform for realizing extraordinary topological phases, attracting considerable attention in recent years Mogi et al. (2017a, b); Xiao et al. (2018); Mogi et al. (2022); Qin et al. (2023a); Chen and Shen (2024); Chang et al. (2013); Yoshimi et al. (2015); Shan et al. (2010); Lu et al. (2010); Li et al. (2010); Lu et al. (2013); Chen et al. (2019); Sun et al. (2020); Liu et al. (2010); Dabiri et al. (2021); Dabiri and Cheraghchi (2021); Sun et al. (2023); Qin et al. (2022a); Zhu et al. (2023). For instance, the well-known Weyl semimetal phase can be interpreted as a stack of 2D Chern insulators with orbital-dependent interlayer tunneling Yang et al. (2011); Chen et al. (2015, 2018a, 2018b); Li et al. (2021a). Similarly, phases such as the axion insulator and the hinged quantum spin Hall effect arise from alternating stacks of Chern insulators with opposite Chern numbers Chen et al. (2021, 2024a, 2023a, 2023b). Moreover, higher-Chern-number quantum anomalous Hall insulators can be constructed by alternating stacks of Chern insulators and normal insulators Zhao et al. (2022); Baba et al. (2022); Li et al. (2022, 2023).
@@
Zhang et al. Zhang et al. (2023) recently proposed and studied a 2D phase modulated by a periodic potential. Unlike conventional Chern insulators, where the spectral flow of edge modes is momentum dependent, their system exhibits edge modes whose spectral flow depends on spatial position. They also suggested that such a periodic potential could be experimentally realized via charge density wave effects Huang et al. (2021); Qin et al. (2020). In a related development, Chen et al. Chen et al. (2024b) explored a 2D higher-order topological insulator and demonstrated that the energies of the corner modes can be tuned by a periodic spin-orbital potential. Additionally, it has been shown that corner modes can also be controlled via long-range hoppings Qin and Chen (2025).
@@
Off-resonant light serves as a powerful tool for inducing Floquet Weyl semimetals Wang et al. (2014); Hübener et al. (2017). Building on Floquet theory Oka and Aoki (2009); Zhan et al. (2024); Qin et al. (2023a, 2022b, 2022a, 2024a); Dabiri et al. (2021); Dabiri and Cheraghchi (2021); Dabiri et al. (2022); Cheraghchi and Askarpour (2023); Pervishko et al. (2018); Calvo et al. (2015); Seshadri and Sen (2022); Seshadri et al. (2019); Seshadri (2023); Lee et al. (2018); Magnus (1954); Blanes et al. (2009); Bukov et al. (2015); Eckardt and Anisimovas (2015); Chen et al. (2018c, a); Du et al. (2022); Wang et al. (2023), it would be intriguing to investigate the experimental implications of realistic periodic driving frequencies. This is particularly relevant in light of recent developments, such as the photoinduced anomalous Hall effect in graphene McIver et al. (2020) and its corresponding microscopic theory Sato et al. (2019), the experimental observation of Floquet states in graphene Merboldt et al. (2025), and the direct detection of Floquet-Bloch states in monolayer graphene Choi et al. (2025), WSe2 Aeschlimann et al. (2021), and the 3D topological insulator Bi2Se3 Wang et al. (2013). Additionally, authors of experimental studies on Floquet transport effects, including anomalous linear transport in light-driven 3D Dirac electrons in bismuth Hirai et al. (2023) and nonperturbative nonlinear transport in a magnetic Floquet-Weyl semimetal Day et al. (2024), have further highlighted the potential of this approach.
@@
In this work, we propose and study a 3D topological phase resembling a Weyl semimetal, modulated by a periodic potential and realized through a simple stack of 2D Chern insulators. The system features Weyl-like points defined in the parameter space $(k_{x},k_{y},z)$, distinct from the conventional Weyl points in momentum space $(k_{x},k_{y},k_{z})$. This Weyl-semimetal-like phase exhibits key properties of a Weyl semimetal, including linear dispersion near the Weyl-like points, nontrivial bulk topology, the emergence of Fermi arc connecting these points, and the Berry monopoles. However, these behaviors are observed in real space rather than momentum space, as in traditional Weyl semimetals. We further compute the local density of states, the layer Hall conductance, and the total 3D Hall conductivity, demonstrating that the Weyl-semimetal-like phase remains robust under weak and moderate interlayer couplings. The impact of disorder is also analyzed: Beyond a critical disorder strength, the Weyl-like points destabilize and the topological phase collapses. In addition, we evaluate the Floquet Chern number under high-frequency driving and show that the positions of the Weyl-like nodes can be effectively tuned by laser fields. Finally, we reveal the emergence of both type-I and II Weyl-like behaviors in a tilted Weyl-like model.
@@
The structure of this paper is as follows. In Section II, we introduce the Weyl-like model Hamiltonian with a periodic potential along the $z$-direction, alongside the conventional Weyl semimetal model, and discuss the duality between these two models. Section III presents calculations of the local density of states, the layer Hall conductances, and the total 3D Hall conductivity in the presence of interlayer hopping for the Weyl-like model. In Section IV, we investigate the effects of a generic disorder potential on the Weyl-like points and the associated topological invariants. In Section V, we derive the Floquet effective Hamiltonians for both models in Section II under high-frequency laser driving. Section VI details the analysis of Floquet Weyl nodes, Floquet Chern numbers, the Floquet energy bands, and Floquet Fermi arcs for both models. A brief discussion of possible experimental realizations of the Weyl-like model is given in Section VII. In Section VIII, we investigate the realization for both type-I and II Weyl-like model behaviors in a tilted Weyl-like model. Finally, Section IX summarizes the main results of the paper.
@@ II Model
In this section, we will introduce the Weyl-like and Weyl semimetal models.
@@ II.1 Weyl-like model
We begin by introducing a two-band model Hamiltonian that describes a stack of 2D Chern insulators subjected to a periodic potential along the $z$ direction. This model is defined in the $(k_{x},k_{y},z)$ space as
${\cal H}({\bf k}_{||},z)$ $\!=\!$ $h_{\text{2D}}({\bf k}_{||})\!+\!V(z)\sigma_{z},$ (1)
where $h_{\text{2D}}({\bf k}_{||})\!=\!m_{z}\sigma_{z}\!+\!m_{0}[2\!-\!\cos(k_{x}a)\!-\!\cos(k_{y}a)]\sigma_{z}\!+\!t_{x}\sin(k_{x}a)\sigma_{x}\!+\!t_{y}\sin(k_{y}a)\sigma_{y}$ represents the Hamiltonian of a 2D Chern insulator when $m_{z}\!\times\!m_{0}\!<\!0$ Shen (2017), and ${\bf k}_{||}\!=\!(k_{x},k_{y})$. Here, $a$ is the lattice constant, and $\sigma_{i}$ (i=x,y,z)(i\!=\!x,y,z) are Pauli matrices acting on the orbital space. The spatially periodic potential is given by $V(z)\!=\!V_{z}\cos(Q_{z}z)$, where $V_{z}$ is the potential strength, $z\!=\!j_{z}a$ denotes the position of the $j_{z}$th layer along the $z$ direction, and $\lambda_{z}\!=\!2\pi/Q_{z}$ is the wavelength of the modulation.
@@
Figure 1: Duality between the decoupled Weyl-like model (left column) and the Weyl semimetal model (right column). The two green vertical dashed lines indicate the positions of the two Weyl-like (or Weyl) points. (a1) Schematic of the decoupled Weyl-like model in real space, featuring a spatially periodic potential $V(z)\!=\!V_{z}\cos(Q_{z}z)$, where $Q_{z}\!=\!2\pi/\lambda_{z}$. (b1) Weyl-like cones in the $(k_{x},k_{y},z)$ space at two critical planes, $(k_{x},k_{y},z_{c-})$ and $(k_{x},k_{y},z_{c+})$, marking topological transitions. (c1) Chern number of each layer [Eq. (3)] as a function of $z$ for the decoupled Weyl-like model with $V_{z}\!=\!1$ eV, $\lambda_{z}\!=\!L_{z}\!=\!31a$, and the lattice constant a=1a\!=\!1 nm. (d1) Energy bands [Eq. (6)] (blue solid curves) of the Weyl-like model at $k_{x}\!=\!k_{y}\!=\!0$. (e1) Energy bands (blue solid curves) and Fermi-like arc (blue horizontal line for E=0E\!=\!0) for the Weyl-like model under open boundary conditions (OBCs) along the $y$ direction, with $k_{x}\!=\!0$ and system size $L_{y}\!=\!31a$. (a2) Schematic of the Weyl semimetal model in real space, featuring a momentum-dependent potential $U(k_{z})\!=\!t_{z}\cos(k_{z}a)$. (b2) Weyl cones in momentum space $(k_{x},k_{y},k_{z})$ at two critical planes $(k_{x},k_{y},k_{c-})$ and $(k_{x},k_{y},k_{c+})$, indicating the topological features of the model. (c2) Chern number [Eq. (II.2)] as a function of $k_{z}$ for the Weyl semimetal model with $t_{z}\!=\!1$ eV. (d2) Energy bands [Eq. (13)] (black solid curves) of the Weyl semimetal model at $k_{x}\!=\!k_{y}\!=\!0$. (e2) Energy bands (black solid curves) and Fermi arc (black horizontal line at E=0E\!=\!0) for the Weyl semimetal model under OBCs along the $y$ direction, with $k_{x}\!=\!0$ and system size $L_{y}\!=\!31a$. Other parameters used: $m_{z}\!=\!0.5$ eV, $m_{0}\!=\!1$ eV, $t_{x}\!=\!t_{y}\!=\!t_{0}\!=\!1$ eV, and $L_{z}\!=\!31a$.
As illustrated in Fig. 1(a1), the system described by the Weyl-like Hamiltonian in Eq. (1) can be interpreted as an array of decoupled 2D layers, each with a spatially dependent effective mass term $\tilde{m}_{z}(z)\!=\!m_{z}\!+\!V(z)$. For the calculations presented in this work, we adopt the fixed parameters $m_{z}\!=\!0.5$ eV, $m_{0}\!=\!1$ eV, $t_{x}\!=\!t_{y}\!=\!t_{0}\!=\!1$ eV, $V_{z}\!=\!1$ eV, and $\lambda_{z}\!=\!31a$, with a=1a\!=\!1 nm. The detailed matrix representation of the real-space Hamiltonian with the periodic potential along the $z$ direction, assuming open boundary conditions (OBCs) along the $y$ direction and periodic boundary conditions (PBCs) along the $x$ direction, is provided in Section SI in the Supplemental Material Sup .
@@
Since the layers are decoupled from each other, we can define a $z$-dependent Chern number of each layer for the $m$th occupied band as Fukui et al. (2005); Sun et al. (2020); Otrokov et al. (2019); Gao et al. (2021); Chen et al. (2024a); Mong et al. (2010); Essin et al. (2009); Wang et al. (2015); Varnava and Vanderbilt (2018); Chen and Shen (2024); Fu et al. (2021); Price et al. (2015); Zhang et al. (2019)
@@
The $z$-dependent Chern number of each layer is illustrated in Fig. 1(c1). Figures 1(a1) and 1(b1) show that each layer behaves as a Chern insulator with Chern number |C|=1|C|\!=\!1 for $z_{c-}\!<\!z\!<\!z_{c+}$ and as a normal insulator with C=0C\!=\!0 outside this region (i.e., for $z\!<\!z_{c-}$ and $z\!>\!z_{c+}$). To determine the critical values $z_{c\pm}$, we evaluate the eigenenergies of the Weyl-like Hamiltonian in Eq. (1) at the high-symmetry point $k_{x}\!=\!k_{y}\!=\!0$, obtaining from Eq. (II.1)
@@
Furthermore, the critical values $z_{c\pm}$ correspond to the gap-closing points, where $E_{\pm}\!=\!\pm\tilde{m}_{z}(z)\!=\!\pm\left[m_{z}\!+\!V_{z}\cos(Q_{z}z)\right]\!=\!0$, leading to
@@
under the condition $1\!\leqslant\!m_{z}/V_{z}\!\leqslant\!1$.
Figure 1(d1) shows the energy band as a function of position $z$ at $k_{x}\!=\!k_{y}\!=\!0$. A clear band inversion occurs near $z\!=\!z_{c\pm}$, highlighting the spectral flow characteristic of the layered topological structure. Additionally, Fig. 1(e1) reveals the presence of a Fermi arc, further demonstrating the topological nature of the system.
Moreover, the total 3D Hall conductivity for the $m$th occupied band is given by Burkov and Balents (2011)
$\sigma_{H}^{3D}$ $\!=\!$ $\sum_{j_{z}=0}^{L_{z}/a}\frac{\sigma_{H}^{2D}(z)}{L_{z}}\!=\!\frac{e^{2}}{h}\!\sum_{j_{z}=0}^{L_{z}/a}\!\frac{C_{m}(z)}{L_{z}},$ (8)
where $j_{z}\!=\!z/a$ is the layer index, $L_{z}$ is the system length along the $z$ direction, $\sigma_{H}^{2D}(z)\!=\!(e^{2}/h)C_{m}(z)$ is the Hall conductance of each layer associated with the $m$th occupied band, $C_{m}(z)$ is defined in Eq. (3), and $L_{z}$ denotes the system size along the $z$ direction. Using the critical values $z_{c\pm}$ obtained in Eq. (7), we can simplify the 3D Hall conductivity expression as
$\sigma_{H}^{3D}$ $\!=\!$ $\frac{e^{2}}{h}\sum_{j_{z}=z_{c-}/a}^{z_{c+}/a}\frac{C_{m}(z)}{L_{z}}$ (9)
$\!\approx\!$ $\!\frac{e^{2}\lambda_{z}}{hL_{z}a\pi}\arccos\left(\frac{m_{z}}{V_{z}}\right)\!,$
@@
The specific choice of the finite system length $L_{z}$ along the $z$ direction plays a crucial role in the topological analysis. In our numerical calculations, we take a finite system length $L_{z}\!=\!\lambda_{z}$ along the $z$ direction, where $\lambda_{z}$ is the modulation period such that $V(z)\!=\!V(z\!+\!\lambda_{z})$. Under this choice, two Weyl-like points appear along the $z$ direction in the finite system.
If a shorter length is chosen, for example, $L_{z}\!<\!z_{c-}$, no Weyl-like point exists in the finite system. If $z_{c-}\!<\!L_{z}\!<\!z_{c+}$, there will be only one Weyl-like point at $z_{c-}$. Conversely, for a longer length, such as $L_{z}\!=\!2\lambda_{z}$, four Weyl-like points will emerge in the finite system.
Therefore, although $V(z)$ is spatially periodic along $z$, the way this periodicity enters the topological analysis depends on the specific choice of the finite system length $L_{z}$ in the $z$ direction.